OpenStudy (anonymous):

If y varies directly as x, and y = 15 as x = -3, find y for the x-value of 4.

OpenStudy (anonymous):

for direct proportionality x1/y1 = x2/y2 -3/15 = 4/y solve

OpenStudy (anonymous):

tell me if u get stuck :D

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

:D